OpenStudy (anonymous):

There are two consecutive even numbers such that five times the smaller minus three times the greater is 6. What is the smaller number?

OpenStudy (dumbcow):

even number is 2n, next even is 2n+2 5(2n) - 3(2n+2) = 6 distribute and add like terms 10n - 6n -6 = 4n -6 = 6 solve for n 4n = 12 n = 12/4 = 3 smaller num is 2n 2n = 6

OpenStudy (anonymous):

If x (the smaller), the greater will be y=(x+2) According to your problem 5x-3(x+2)=6 5x-3x-6=6 2x=12 2x=12 x=12/2 x=6

OpenStudy (anonymous):

Do you have the answer?

OpenStudy (anonymous):

become my fan please